# RUSplatting: Robust 3D Gaussian Splatting for Sparse-View Underwater Scene Reconstruction

### v1

@@ 5 Experiments
Instant-NGP [22] 23.7944 0.6317 0.4434 18.9001 0.4693 0.3966 19.0266 0.5107 0.4977
SeaThru-NeRF [17] 26.0604 0.7806 0.3398 21.2253 0.5948 0.4104 19.1967 0.4967 0.5685
3DGS [15] 26.8353 0.9011 0.1826 23.7563 0.7745 0.2231 22.6551 0.7544 0.3288
WaterSplatting [18] 27.3924 0.8475 0.1870 24.8788 0.7915 0.2159 24.5832 0.7517 0.3172
UW-GS [27] 28.2557 0.9105 0.1750 24.1833 0.7838 0.2050 24.3409 0.7723 0.3107
RUSplatting (ours) 29.2928 0.9163 0.1665 25.5955 0.8138 0.1964 25.7990 0.7724 0.3008
@@ 5.1 Implementation details
Three datasets were used: SeaThru-NeRF [17], S-UW [27] and our newly collected dataset, Submerged3D. Submerged3D consists of four scenes, each containing 20 RGB 720p images focusing on shipwrecks captured in the deep sea. It is challenging for 3D reconstruction due to sparse viewpoints and low-light environment conditions. More details about the objects and sites, as well as the public link to the dataset, are omitted for double-blind review and will be included in the final accepted version. Example images can be found in the supplementary material. To create the training and test splits, we assign every 8th image in the dataset to the test set and use the rest for training.
@@
We utilised COLMAP [25] to initialise the point cloud and estimate the camera poses for the image sequences. For each pair of adjacent images, only one interpolated frame is generated. The experiments are conducted on a single RTX 4090 GPU. All experiments are run for 20,000 iterations. $\lambda_{r}$, $\lambda_{d}$, and $\lambda_{ca}$ are set to 0.8, 0.1, and 1, respectively. The detailed datasets can be found at here, and the implementation of this project is available at repository.
@@ 5.2 Evaluation
We compared our results with the Instant-NGP [22], SeaThru-NeRF [17], 3DGS [15], WaterSplatting [18] and UW-GS [27]. For a fair comparison, we used their official implementations and trained each method on the same image sequences using our dataset split strategy. The quality of the synthesised views was evaluated using PSNR, SSIM, and LPIPS metrics.
@@ 6 Conclusion
This paper presents a new GS-based framework for deep underwater scene reconstruction. Specifically, we propose decoupled learning for the RGB channels, leveraging wavelength-dependent attenuation and scattering properties of underwater light. This approach enables more accurate colour reconstruction. To mitigate challenges posed by sparse viewpoints, we incorporate a frame interpolation mechanism that facilitates denser and more robust scene capture, along with adaptive frame weighting during training. Finally, we propose a new loss function that dynamically adjusts to image characteristics, effectively suppressing noise while preserving geometric structure. Collectively, our RUSplatting framework significantly enhances the visual fidelity, structural consistency, and colour accuracy of underwater reconstructions.

### v2

@@ 5 Experiments
Instant-NGP [Müller et al.(2022)Müller, Evans, Schied, and Keller] 23.7944 0.6317 0.4434 18.9001 0.4693 0.3966 19.0266 0.5107 0.4977
SeaThru-NeRF [Levy et al.(2023)Levy, Peleg, Pearl, Rosenbaum, Akkaynak, Korman, and Treibitz] 26.0604 0.7806 0.3398 21.2253 0.5948 0.4104 19.1967 0.4967 0.5685
3DGS [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] 26.8353 0.9011 0.1826 23.7563 0.7745 0.2231 22.6551 0.7544 0.3288
[…]
@@ 5.1 Implementation details
Figure 2: Sample images from Submerged3D dataset.
Three datasets were used: SeaThru-NeRF [Levy et al.(2023)Levy, Peleg, Pearl, Rosenbaum, Akkaynak, Korman, and Treibitz], S-UW [Wang et al.(2025)Wang, Anantrasirichai, Zhang, and Bull] and our newly collected dataset, Submerged3D. Submerged3D consists of four scenes, each containing 20 RGB 720p images focusing on shipwrecks captured in the deep sea. It is challenging for 3D reconstruction due to sparse viewpoints and low-light environment conditions. Examples of the scenes in Submerged3D are shown in Fig. 2. To create the training and test splits, we assign every 8th image in the dataset to the test set and use the rest for training. SeaThru-NeRF, S-UW, and our Submerged3D provide no ground-truth depth, therefore we use Depth-Anything-V2 [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] pseudo-depth as weak supervision.
@@
We utilised COLMAP [Schönberger and Frahm(2016)] to initialise the point cloud and estimate the camera poses for the image sequences. For each pair of adjacent images, only one interpolated frame is generated. The experiments are conducted on a single RTX 4090 GPU. All experiments are run for 20,000 iterations. $\lambda_{r}$, $\lambda_{d}$, and $\lambda_{ca}$ are set to 0.8, 0.1, and 1, respectively.
@@ 5.2 Evaluation
We compared our results with the Instant-NGP [Müller et al.(2022)Müller, Evans, Schied, and Keller], SeaThru-NeRF [Levy et al.(2023)Levy, Peleg, Pearl, Rosenbaum, Akkaynak, Korman, and Treibitz], 3DGS [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis], WaterSplatting [Li et al.(2025)Li, Song, Xu, Elsig, and Kulhanek] and UW-GS [Wang et al.(2025)Wang, Anantrasirichai, Zhang, and Bull]. For a fair comparison, we used their official implementations and trained each method on the same image sequences using our dataset split strategy. The quality of the synthesised views was evaluated using PSNR, SSIM, and LPIPS metrics.
@@ 6 Conclusion
This paper presents a new GS-based framework for deep underwater scene reconstruction. Specifically, we propose decoupled learning for the RGB channels, leveraging wavelength-dependent attenuation and scattering properties of underwater light. This approach enables more accurate colour reconstruction. To mitigate challenges posed by sparse viewpoints, we incorporate a frame interpolation mechanism that facilitates denser and more robust scene capture, along with adaptive frame weighting during training. Finally, we propose a new loss function that dynamically adjusts to image characteristics, effectively suppressing noise while preserving geometric structure. Collectively, our RUSplatting framework significantly enhances the visual fidelity, structural consistency, and colour accuracy of underwater reconstructions. However, as RIFE and Depth-Anything-V2 are not specifically designed for underwater conditions, future work could improve their robustness through fine-tuning or physics-based priors, thereby further enhancing the colour fidelity and geometric accuracy of our framework.
Acknowledgements
This work was supported by the UKRI MyWorld Strength in Places Programme(SIPF00006/1) and EPSRC ECR(EP/Y002490/1).